OpenStudy (anonymous):

please help medals will be given how many solutions does the system of equations have? x-14y=12 and 5x -20y=60 one two infinitely many none

OpenStudy (haseeb96):

two

OpenStudy (anonymous):

how many solutions does the system of equations have? y-5x=-6 and 3y-15x =-12 @Haseeb96 what about this one

OpenStudy (anonymous):

Those are two parallel lines that never intersect.. so none

OpenStudy (anonymous):

@ECE th second one i just posted right ?

OpenStudy (haseeb96):

there are two methods by which you can solve this system of equation 1. Elimination method 2. substitution method

OpenStudy (anonymous):

so the answer is ?

OpenStudy (radar):

Using the elimination method, I believe you will find there is only one solution that will satisfy the first system. (12,0)

OpenStudy (radar):

The elimination method is easy enough, Multiply first equation thru by 5, then subtract and solve for y.

OpenStudy (haseeb96):

two solution

OpenStudy (radar):

@Haseeb96 other than x=12, and y = 0, what is another solution for the First system? x-14y=12 and 5x -20y=60

OpenStudy (haseeb96):

that is substitution method in which any one equation you take separate only one variable from the equaation then put in the another equation

OpenStudy (radar):

Are you saying there is two methods to arrive at two different solutions or arrive at a single solution?????

OpenStudy (radar):

Using substitution method: x - 14y = 12 x=14y + 12 Substitute in the 2nd equation 5x - 20y = 60 5(14y + 12) - 20y = 60 70y + 60 - 20y=60 50y = 0 y=0, substiution of 0 for y in either of the original equations result in a value of 12 for x The same solution as obtained by the elimination method.